RECENT DEVELOPMENT IN SUM FREQUENCY GENERATION SPECTROSCOPY SYSTEMS

\underline{MICHAEL PATTERSON}, LUCIAN HAND, ALTOS PHOTONICS, BOZEMAN, MT; ZENONAS KUPRIONAS, UAB EKSPLA, VILNIUS (LITHUANIA).

For more than 20 years, EKSPLA has been pursuing improved picosecond lasers, and related instruments. Working together with SFG researchers around the world, we have introduced a number of innovations which improve this important tool for surface chemistry. Recent advancements include a 50 Hz DPSS pump pump laser, utilization of CAN-USB interface for improved data collection, and a SFG-Microscopy system demonstrated and now being refined. As we move forward, we solicit your feedback and suggestions for improving the functionality and accessibility of SFG Spectroscopy systems.